The amount of P10,000 is to be invested for 1 year. A part of the 10,000 is invested at 6% and the remainder at 7%. The income from the P10,000 is P640 per year. Find the amount invested at each rate.
Answer by macston(5194)

.
X=amount at 6%; Y=amount at 7%
.
X+Y=P10,000
X=P10,000-Y
.
0.06X+0.07Y=P640
0.06(P10000-Y)+0.07Y=P640
P600-0.06Y+0.07Y=P640
0.01Y=P40
Y=P4000 ANSWER 1: P4000 was invested at 7%.
.
X=P10000-Y=P10000-P4000=P6000 ANSWER 2: P6000 was invested at 6%.
.
CHECK:
0.06X+0.07Y=P640
0.06(P6000)+0.07(P4000)=P640
P360+P280=P640
P640=P640